Solve the following system :
-9x + 3y -2z = 61
8x + 7y + 5z = -138
5x - 5y + 8z = -45

Answer by Alan3354(69443)

Solve the following system :
-9x + 3y -2z = 61
8x + 7y + 5z = -138
5x - 5y + 8z = -45
------------------------------------
-36x + 12y -8z = 244 --- 1st eqn times 4
 5x - 5y + 8z = -45
---------------------------- Add
-31x + 7y = 199 ---- z eliminated  (Eqn A)
===================
-45x + 15y - 10z = 305 --- 1st eqn times 5
 16x + 14y + 10z = -276 --- 2nd eqn times 2
-------------------------------------------------- Add
-29x + 29y = 29
 -x + y = 1 --- z eliminated (Eqn B)
============================
-31x + 7y = 199
 -7x + 7y = 7  Eqn B times 7
----------------------------------------
Subtract
-24x = 192
x = -8
---------------
Sub for x in Eqn B find y
Sub x & y in any eqn and find z